OpenStudy (ledah):

After isolating the absolute value in the equation -2|5n + 10| = -30, you would see that there is “no solution” for this problem.

OpenStudy (anonymous):

\[\left| 5n+10 \right|=15 \]

OpenStudy (anonymous):

5n+10=15 and 5n+10=-15

OpenStudy (ledah):

So its true. no solution

OpenStudy (anonymous):

5n=5 n=1 and 5n=-25 n=-5

OpenStudy (anonymous):

how no solution?

OpenStudy (ledah):

actually I confused myself. Hold on...

OpenStudy (ledah):

If what you said is correct. (not doubting) then the solution doesn't quite match up to it correctly.